Find the value of t.

The midpoint of the segment joining (t,2) and (3t, -4) is (-6,-1).

Answer by checkley71(8405)

LENGTH OF THE X LINE IS 3T+T=4T
4T/2=-6 CROSS MULTIPLY.
4T=-12
T=-12/4
T=-3 OR THE POINT (-3,2)
3T=-3*3=-9 OR THE POINT (-9,-4